# P-MSDiff: Parallel Multi-Scale Diffusion for Remote Sensing Image Segmentation

### v1

@@ I Introduction
Remote sensing image segmentation is a critical research area within remote sensing image processing. Its primary objective is to divide remote sensing imagery into separate semantic regions. The utilization of high-resolution or hyperspectral images for precise delineation provides essential spatial information for various applications such as geological exploration, environmental monitoring, and urban planning. The utilization and implementation of high-resolution remote sensing data are becoming more demanding in terms of accuracy and efficiency due to the progress in satellite and drone technologies. In response, researchers have devised a variety of deep learning-based semantic segmentation algorithms tailored for remote sensing imagery. The effective integration and swift advancement of deep convolutional neural networks (CNNs)[1, 2, 3] have resulted in remarkable performance across a range of tasks, such as image classification[4, 5], image segmentation[6, 7], instance segmentation[8, 9], image vectorization[10], and object detection[11, 12, 13].
@@ IV-B1 Comparative Methods
Our model consists of several key components: The Time Step Encoding Module, which documents the correlation between diffusion times and model weights, and then incorporates this information into the encoding phase; the CNN-Transformer Encoder Architecture, which operates simultaneously during both the image encoding and segmentation mask encoding phases; the Decoding Process, which is closely integrated with the encoding phase, progressively upsamples while incorporating lower-dimensional features from the encoder of the same resolution; the Cross-Level Bilateral Attention (CBLA) Module, which is embedded in each downsampling stage of the encoder and each upsampling stage of the decoder; Sub-branches, with fewer structural layers, reconstruct the core of the encoder-decoder architecture, merging with the main framework at different points in the decoding process.

### v2

@@ I Introduction
Remote sensing image segmentation is a critical research area within remote sensing image processing. Its primary objective is to divide remote sensing imagery into separate semantic regions. The utilization of high-resolution or hyperspectral images for precise delineation provides essential spatial information for various applications such as geological exploration, environmental monitoring, and urban planning. The utilization and implementation of high-resolution remote sensing data are becoming more demanding in terms of accuracy and efficiency due to the progress in satellite and drone technologies. In response, researchers have devised a variety of deep learning-based semantic segmentation algorithms tailored for remote sensing imagery. The effective integration and swift advancement of deep convolutional neural networks (CNNs)[1, 2, 3] have resulted in remarkable performance across a range of tasks, such as image classification[4, 5], image segmentation[6, 7], instance segmentation[8, RN09], image vectorization[10], and object detection[11, 12, 13].
@@ IV-B1 Comparative Methods
Our model consists of several key components: The Time Step Encoding Module, which documents the correlation between diffusion times and model weights, and then incorporates this information into the encoding phase; the CNN-Transformer Encoder Architecture, which operates simultaneously during both the image encoding and segmentation mask encoding phases; the Decoding Process, which is closely integrated with the encoding phase, progressively upsamples while incorporating lower-dimensional features from the encoder of the same resolution; the Cross-Bridge Linear Attention (CBLA) Module, which is embedded in each downsampling stage of the encoder and each upsampling stage of the decoder; Sub-branches, with fewer structural layers, reconstruct the core of the encoder-decoder architecture, merging with the main framework at different points in the decoding process.